write liner equations in slope-intercept form for each of the following:
1. Through(2,6) and m= 2/3
2 Through (-3,5) and parallel to y=-4x+2
3. Horizontal line through (-9,3)
4. through(2,-7) and (-1,-10)

Answer by richwmiller(17219)

#1
the slope intercept form
is y=mx+b
where m is the slope and b is the y intercept
first take y=mx+b and plug in the x and y and solve for b
then take y=mx+b and put in the b that you found and the m that was given.
#2lines that are parallel have the same slope
#3horizontal lines have a slope of 0
where will the y intercept be. if y is 0 and x is always the same?
#4to find the slope knowing two points
take the difference of the y's (subtract them) and divide that by the difference of the x's
that is the slope
then do the same as we did for the first problem